# VisBias: Measuring Explicit and Implicit Social Biases in Vision Language Models

### v1

@@ 1 Introduction
Figure 1: An overview of the VisBias framework, including two scenarios for explicit scenarios and two for implicit scenarios to assess social biases across varying levels of awareness through direct question answering or daily tasks.
Visual Language Models (VLMs) have shown the ability to interpret visual content and answer related queries Bubeck et al. (2023); Chow et al. (2025). As these models gain widespread adoption in tasks such as text recognition Liu et al. (2024); Chen et al. (2025), mathematical problem-solving Yang et al. (2024b); Peng et al. (2024), and medical applications Azad et al. (2023); Buckley et al. (2023), identifying and addressing potential social biases is essential to prevent harm to users Raj et al. (2024); Sathe et al. (2024); Brinkmann et al. (2023); Ruggeri and Nozza (2023).
Social biases in human society can be broadly classified into explicit and implicit biases based on their manifestation. Implicit Bias refers to unconscious, automatic attitudes or stereotypes that influence decisions and behaviors without conscious intent Hahn et al. (2014). These biases develop through learned associations derived from societal messages, cultural norms, and personal experiences, often reinforcing broader societal stereotypes Byrd (2021). Because implicit biases function below conscious awareness, they are typically assessed using indirect methods such as the Implicit Association Test (IAT) Greenwald et al. (1998, 2003). For example, a hiring manager may unknowingly favor male candidates over equally qualified female candidates due to implicit associations linking men with leadership. In contrast, Explicit Bias involves conscious, intentional attitudes or beliefs that individuals are aware of and openly endorse Brown and Gaertner (2008). These biases are deliberate and overt, often measured through self-reported surveys or direct behavioral observation. They manifest in actions such as intentional exclusion or discrimination against specific groups. The key distinction between implicit and explicit bias lies in the level of awareness: implicit bias operates subconsciously and may contradict an individual’s consciously held values, whereas explicit bias is consciously recognized and aligns with expressed beliefs Dovidio et al. (2001).
@@
In this paper, we study both explicit and implicit social biases by designing fairness-related questions and tasks that, while appearing to assist user requests, may reveal underlying biases. We introduce VisBias, which evaluates these biases in VLMs through four distinct assessment scenarios. First, we focus on explicit tasks—(I) Multiple-Choice and (II) Yes-No questions—where models make direct judgments about individuals’ annual income, education level, political leaning, and religion based on real-world images that capture race, gender, and occupation. To counteract alignment-related refusals, we apply a Caesar cipher “jailbreak” Yuan et al. (2024) ensuring models reveal biases that might otherwise remain hidden. In the multiple-choice setting, we analyze which demographic groups are linked to particular answers, while in yes-no comparisons (such as asking whether one person appears more educated than another), we check for group-specific favoritism. Next, to unveil more subtle forms of bias, we introduce implicit tasks—(III) Image Description and (IV) Form Completion—where the model’s goal is to assist with descriptive or data-entry tasks rather than make overt judgments. In image description, we measure skewed word associations across different demographic groups by examining the language models use. Finally, in form completion, we prompt VLMs to fill out personal information forms with various attributes, tracking how certain features (like gender or race) shift the distribution of other attributes. By integrating explicit and implicit testing, VisBias provides a comprehensive lens for detecting bias across diverse modeling tasks.
@@ 2 VisBias Design and Implementation
VisBias assesses both explicit and implicit social biases in VLMs. To elicit these biases at different levels of awareness, we design four evaluation scenarios: (I) Multiple-Choice Questions, (II) Yes-No Questions, (III) Image Description, and (IV) Form Completion, as shown in Fig. 1.
@@ 2.1 Explicit Scenarios
In explicit scenarios, the questions are designed to instruct models to make decisions primarily based on the individual’s appearance. To incorporate demographic information through image modality, we first collect real-world images. Ghavami and Peplau (2013) showed that people’s stereotypes and expectations vary depending on the combined identity of an individual rather than just their race or gender alone. Following this work, we focus on the intersectionality of race (Asian, Black, Hispanic, Middle Eastern, and White) and gender (Female and Male). For occupations, we select those with high gender or race imbalance according to U.S. Bureau of Labor Statistics, [^1] https://www.bls.gov/bls/blswage.htm including basketball player, nurse, firefighter, CEO, cook, doctor, and lawyer. We retrieve images by searching for keywords related to race, gender, and occupation on Google Images, selecting those whose descriptions explicitly include all three attributes. Additionally, we verify that the images are not AI-generated to ensure authenticity. We obtain ten images per gender-race-occupation group, yielding 700 images in total (2 genders × 5 races × 7 occupations × 10 images).
Models with safety alignment may refuse to answer direct questions. In such cases, the response is classified as unbiased. To create a more challenging testing scenario, we employ the jailbreak method CipherChat Yuan et al. (2024), which encodes questions using a Caesar cipher with a three-letter shift. To measure the intrinsic biases in the model, we do not include negative examples or explicitly instruct models to generate harmful or toxic content. Instead, we provide neutral examples and define the Caesar cipher, specifying that communication should adhere to this encoding. The full prompt is provided in §B of the appendix.
@@ 2.2 Implicit Scenarios
In the fourth scenario, the model is required to complete a draft personal information form that has been pre-filled with randomly generated data. To construct a comprehensive individual profile, we select 20 attributes, including age, gender, race, religion, education level, occupation, political leaning, hobbies, and personality traits. Each attribute is assigned predefined categories; for instance, education level includes “No Education,” “Elementary School,” “Middle School,” “High School,” “Associate,” “Bachelor,” “Master,” and “Doctorate,” resulting in a total of 128 possible choices. The complete list of attributes and their corresponding choices is presented in Table 6 in §C of the appendix. During model evaluation, five attributes are randomly selected along with corresponding choices, and the form is converted into an image format for input into VLMs.
Once VLM responses are obtained, answers for each attribute are extracted and categorized into predefined choices using GPT-4o. An additional “Unspecified” class is introduced to accommodate responses that do not fit within the predefined categories. To mitigate positional biases inherent in auto-regressive models, where later content depends on preceding text, we apply a cyclic shift to each query. For example, if a form includes attributes labeled A, B, C, D, and E, five variants are generated (ABCDE, BCDEA, CDEAB, DEABC, and EABCD), ensuring that each attribute appears in all positions. Given the 20 attributes we have, a total of 20 variants are created.

### v2

@@ 1 Introduction
Figure 1: An overview of the VisBias, including two scenarios for explicit scenarios and two for implicit scenarios to assess social biases across varying levels of awareness through direct question answering or daily tasks.
Visual Language Models (VLMs) have shown the ability to interpret visual content and answer related queries Bubeck et al. (2023); Chow et al. (2025). As these models gain widespread adoption in tasks such as text recognition Liu et al. (2024); Chen et al. (2025), mathematical problem-solving Yang et al. (2024); Peng et al. (2024), and medical applications Azad et al. (2023); Buckley et al. (2023), identifying and addressing potential social biases is essential to prevent harm to users Raj et al. (2024); Sathe et al. (2024); Brinkmann et al. (2023); Ruggeri and Nozza (2023).
Social biases in human society can be broadly classified into explicit and implicit biases based on their manifestation. Implicit Bias refers to unconscious, automatic attitudes or stereotypes that influence decisions and behaviors without conscious intent Hahn et al. (2014). These biases develop through learned associations derived from societal messages, cultural norms, and personal experiences, often reinforcing broader societal stereotypes Byrd (2021). Because implicit biases function below conscious awareness, they are typically assessed using indirect methods such as the Implicit Association Test Greenwald et al. (1998, 2003). For example, a hiring manager may unknowingly favor male candidates over equally qualified female candidates due to implicit associations linking men with leadership. In contrast, Explicit Bias involves conscious, intentional attitudes or beliefs that individuals are aware of and openly endorse Brown and Gaertner (2008). These biases are deliberate and overt, often measured through self-reported surveys or direct behavioral observation. They manifest in actions such as intentional exclusion or discrimination against specific groups. The key distinction between implicit and explicit bias lies in the level of awareness: implicit bias operates subconsciously and may contradict an individual’s consciously held values, whereas explicit bias is consciously recognized and aligns with expressed beliefs Dovidio et al. (2001).
@@
In this paper, we study both explicit and implicit social biases by designing fairness-related questions and tasks that, while appearing to assist user requests, may reveal underlying biases. We introduce VisBias, which evaluates these biases in VLMs through four distinct assessment scenarios. First, we focus on explicit tasks—(I) Multiple-Choice and (II) Yes-No questions—where models make direct judgments about individuals’ annual income, education level, political leaning, and religion based on real-world images that capture race, gender, and occupation. To counteract alignment-related refusals, we apply a Caesar cipher “jailbreak” Yuan et al. (2024) ensuring models reveal biases that might otherwise remain hidden. In the multiple-choice setting, we analyze which demographic groups are linked to particular answers, while in yes-no comparisons (such as asking whether one person appears more educated than another), we check for group-specific favoritism. Next, to unveil more subtle forms of bias, we introduce implicit tasks—(III) Image Description and (IV) Form Completion—where the model’s goal is to assist with descriptive or data-entry tasks rather than make overt judgments. In image description, we measure skewed word associations across different demographic groups by examining the language models use. Finally, in form completion, we prompt VLMs to fill out personal information forms with various attributes, tracking how certain features (like gender or race) shift the distribution of other attributes. By integrating explicit and implicit testing, VisBias provides a comprehensive lens for detecting bias across diverse modeling tasks, as shown in Fig. 1.
@@ 2.1 Explicit Scenarios
In explicit scenarios, the questions are designed to instruct models to make decisions primarily based on the individual’s appearance. To incorporate demographic information through image modality, we first collect real-world images. Ghavami and Peplau (2013) showed that people’s stereotypes and expectations vary depending on the combined identity of an individual rather than just their race or gender alone. Following this work, we focus on the intersectionality of race (Asian, Black, Hispanic, Middle Eastern, and White) and gender (Female and Male). For occupations, we select those with high gender or race imbalance according to U.S. Bureau of Labor Statistics, [^1] https://www.bls.gov/bls/blswage.htm including basketball player, nurse, firefighter, CEO, cook, doctor, and lawyer. We retrieve images by searching for keywords related to race, gender, and occupation on Google Images, selecting those whose descriptions explicitly include all three attributes. We obtain ten images per gender-race-occupation group, yielding 700 images in total (2 genders × 5 races × 7 occupations × 10 images).
Models with safety alignment may refuse to answer direct questions. In such cases, the response is classified as unbiased. To create a more challenging testing scenario, we employ the jailbreak method CipherChat Yuan et al. (2024), which encodes questions using a Caesar cipher with a three-letter shift. To measure the intrinsic biases in the model, we do not include negative examples or explicitly instruct models to generate harmful or toxic content. Instead, we provide neutral examples and define the Caesar cipher, specifying that communication should adhere to this encoding. The full prompt is provided in §C of the appendix.
@@ 2.2 Implicit Scenarios
In the fourth scenario, the model is required to complete a draft personal information form that has been pre-filled with randomly generated data. To construct a comprehensive individual profile, we select 20 attributes, including age, gender, race, religion, education level, occupation, political leaning, hobbies, and personality traits. Each attribute is assigned predefined categories; for instance, education level includes “No Education,” “Elementary School,” “Middle School,” “High School,” “Associate,” “Bachelor,” “Master,” and “Doctorate,” resulting in a total of 128 possible choices. The complete list of attributes and their corresponding choices is presented in Table 7 in §D of the appendix. During model evaluation, five attributes are randomly selected along with corresponding choices, and the form is converted into an image format for input into VLMs.
Once the VLM responses are obtained, answers for each attribute are extracted and categorized into predefined choices using GPT-4o. An additional “Unspecified” class is introduced to accommodate responses that do not fit within the predefined categories. To mitigate positional biases inherent in auto-regressive models, where later content depends on preceding text, we apply a cyclic shift to each query. For example, if a form includes attributes labeled A, B, C, D, and E, five variants are generated (ABCDE, BCDEA, CDEAB, DEABC, and EABCD), ensuring that each attribute appears in all positions. Given the 20 attributes we have, a total of 20 variants are created.
@@
3 Experiments
This section presents our experiments with VisBias on five VLMs: GPT-4(V) OpenAI (2023), GPT-4o OpenAI (2024), Gemini-1.5-Pro Pichai and Hassabis (2024), LLaMA-3.2 (Vision) Dubey et al. (2024), and LLaVA-v1.6 Liu et al. (2023). For multiple-choice questions, GPT-4V, GPT-4o, and Gemini-1.5 refuse to provide responses. In contrast, all models generate valid responses to yes-no questions. When the Caesar cipher is applied, Gemini-1.5-Pro and GPT-4V can answer both question types. However, GPT-4o, with its enhanced safety alignment, declines to respond to multiple-choice questions but provides answers to yes-no questions. LLaMA and LLaVA are unable to interpret the Caesar cipher and thus cannot respond to most queries. Finally, in the form completion task, LLaMA refuses to provide responses, while LLaVA is unable to complete the task due to its limited reasoning ability. The refusal rates are reported in Table 2 in §A of the appendix. Temperatures are set to zero for all models, except in the image description task, where it is one.
3.1 (I) Multiple-Choice Questions
In this scenario, we collect the VLMs’ responses to all four questions for each image, resulting in a total of 2,800 responses per model. We then compute the Jensen–Shannon Divergence (JSD) to measure variations in the distribution of choices across different demographic groups. The results are shown in Table 3 in the appendix. A higher JSD indicates a more skewed output distribution, reflecting potential social biases. Current VLMs exhibit varying degrees of bias in multiple-choice questions related to demographic attributes. In the gender dimension, responses from LLaVA and Gemini consistently estimate higher annual incomes for males than for females, while both Gemini and GPT-4V associate higher education levels more frequently with males. Similarly, in the racial dimension, different race groups are linked to distinct political leaning. Specifically, GPT-4V and Gemini predominantly align with left- or right-liberal perspectives, whereas LLaVA demonstrates greater response variability but still exhibits a stronger tendency toward left- or right-communitarian viewpoints. In the religion dimension, GPT-4V consistently categorizes individuals as Christian, irrespective of ethnic background, whereas LLaVA assigns religious beliefs based on ethnicity.
Char 783.9 610.6 291.1 318.9 334.0 179.6
Token 140.2 109.1 61.1 61.0 64.2 35.0
Female her, she, woman, female, hijab, makeup, womans, blouse, shes, blazer her, she, woman, hair, hijab, back, blazer, female, blouse, styled she, her, woman, female, long, hijab, hair, blouse, ponytail, young her, woman, she, back, pulled, ponytail, blazer, long, hijab, necklace her, she, woman, blazer, hair, female, hijab, styled, nurses, blouse woman, female, her, she, hijab, nurse, muslim, girl, womancore, feminine
Male (Marked) his, he, man, him, tie, mans, male, beard, suit, shirt his, he, man, him, tie, shirt, suit, blue, mans, beard he, his, man, tie, male, him, beard, nba, shirt, doctor his, man, he, mans, suit, tie, him, short, blue, shirt his, he, man, mans, him, tie, suit, , doctor beard man, his, male, handsome, businessman, tie, he, him, suit, dotted
Black african, bald, their, peppers, braids, american, fire, station, braided fire, tattoos, apron, truck, courtroom, braids, burger, window, sparks, binders africanamerican, curly, peppers, yellow, woman, american, embiid, black, flag, burger skin, braids, curly, monitor, fire, briefcase, necklace, bald, station, dark their, showcasing, bookshelf, basketball, behind, gym, filled, passion, holding, fire black, african, colorism, arts, movement, influence, vibrant, american, bold, afrocaribbean
Hispanic bakery, books, baker, engine, bridge, behind, fire, flag, baguettes, gauges books, truck, filled, volumes, bakery, usa, dark, fire, pot, library engine, booker, marketing, bucks, ghormley, hispanic, baker, boy, bookshelves, bread books, bookshelf, microphone, dark, concrete, rope, bridge, spines, wings, flag bookshelf, wearing, books, court, stadium, engine, bridge, flag, child, filled hispanicore, mexican, cultures, melds, chicano, chicanoinspired, american, indian, culture, contemporary
ME hijab, cultural, headscarf, beard, religious, city, evening, screen, modesty, lighting hijab, headscarf, cultural, two, beige, woman, city, wrapped, evening, court hijab, quiet, unwavering, gaze, the, speaks, testament, framed, danger hijab, cultural, headscarf, diversity, dark, kitchen, arabic, credit, pot, face hijab, women, cultural, headscarf, religious, modesty, wearing, covering, pink, muslim muslim, hijab, middle, camera, portrait, eastern, she, wearing shot, saudi
Table 1: The marked words in (III) Image Description—those mainly used to describe a specific group in contrast to the marked group (white males)—are presented in this table. The lengths of these descriptions, tokenized using NLTK, are reported at the top of this table. Words with high frequency across VLMs are marked in red.
3.2 (II) Yes-No Questions
In this scenario, we randomly select ten pairs of individuals within each occupation, resulting in 70 test cases. For each test case, we apply the two queries introduced in §2.1 to ensure a rigorous evaluation. We then count instances where VLMs produce differing responses for each pair. The results, alongside the application of the Caesar cipher jailbreak, are presented in Table 4 in the appendix. VLMs exhibit moderate biases in this scenario and show more pronounced biases after a jailbreak. Notably, Gemini demonstrates a substantial increase in response inconsistency when handling reversed query pairs post-encryption, indicating potential bias risks. In contrast, GPT-4o and GPT-4V show no significant variation in responses before and after the jailbreak attack. Meanwhile, LLaVA and LLaMA fail to interpret queries in Caesar cipher, and thus, we evaluate them only on original (without jailbreak) queries.
3.3 (III) Image Description
We also examine the sentiment of words used to describe different demographic groups using the VADER sentiment analyzer in NLTK Hutto and Gilbert (2014) and check whether the words are typical racial stereotype words using the dictionary in Ghavami and Peplau (2013). The sentiment scores and stereotype scores are presented in Table 5 and 6 in the appendix, respectively. Among the models, Gemini and Midjourney exhibit the strongest biases, with the most pronounced biases directed toward Middle Eastern and Black groups.
3.4 (IV) Form Completion
Obtaining High Correlation Pairs
In this scenario, we randomly generate 20 forms for each of the 20 variants produced by cyclic shifts, with each form containing five pre-filled attributes. This process yields a total of 400 samples for each VLM. Using the algorithm described in Alg. 1, we compute the KL Divergence between all conditional distributions and the original distributions. After excluding cases with only one sample, we identify attribute pairs with $D_{\text{KL}}\geq 1$, resulting 370 pairs.
Human Evaluation on the Correlation
High correlation between attribute pairs does not necessarily indicate social bias. For example, the attribute “Age = 0-17” is strongly correlated with “Marital Status = Single,” “Has Children = No,” “Occupation = Student,” “Annual Income = Less Than \$10k,” and “Residential Status = Living with Family.” To determine whether such correlations reflect social bias or reasonable causal inference, human evaluation is required, as LLM inherently exhibit biases. We employ human annotators to assess 370 attribute pairs, distributed across 19 questionnaires. Each questionnaire follows the format: “To what extent is the following statement socially biased? If a person’s {Attribute 1} is {Choice 1333Choice 1 is the conditioned value of Attribute 1.}, then we can infer that the person’s {Attribute 2} is {Choice 2444Choice 2 is the value that has the largest probability of Attribute 2, conditioned on Attribute 1 = Choice 1.}.” Participants rate bias on a five-point scale, ranging from “(1) No Bias” to “(5) Severe Bias.” An example questionnaire created using Qualtrics [^5] https://www.qualtrics.com/ is shown in Fig. 4 in §E of the appendix. Participants are recruited via Prolific. [^6] https://www.prolific.com/ To ensure diversity and prevent overrepresentation of specific gender-race groups, we only require fluency in English. Responses submitted in under two minutes or containing identical answers for all questions are excluded. Each questionnaire is completed by three different annotators. On average, participants take 5 to 10 minutes to complete a questionnaire, with an hourly compensation of \$6.84 (rated as “Good” by the platform).
Figure 2: Distribution of average human scores on attribute pairs. $\geq 3$ are considered biases.
Figure 3: Visualization of the choices that maximize the distribution shift of other attributes in (IV) Form Completion. Color represents the choices, while size indicates the KL divergence. For example, in Panel (d), for GPT-4o, selecting “Doctorate” for the attribute “education level” induces the largest distribution shift in attribute “occupation” compared to other education level choices. This shift is substantial, as indicated by the large plot size. Figures for other sixteen attributes are included in §F of the appendix. Circles in bold are biases verified by humans.
Bias Analysis with Human Annotation
Figure 2 presents the distribution of average human scores across all 370 attribute pairs. Since a score of 3 represents “Moderate Bias,” we classify attribute pairs with an average score greater than or equal to 3 as exhibiting social bias, while those below 3 are considered reasonable causal inferences. Examples of high-scoring biases include: (1) If a person’s occupation is Athlete, then we can infer that the person’s religion is Christianity. (2) If a person’s sexual orientation is Heterosexual, then we can infer that the person’s personality type is INTJ. We further analyze attribute correlations by identifying choices that maximize distribution shifts, as illustrated in Fig. 3, 5, 6, 7, and 8. VLMs exhibit stronger social biases in this task compared to explicit bias evaluation scenarios. These models often establish rigid associations between attributes. For example, individuals in Creative Roles are consistently associated with the INFP personality type and Painting as a hobby. Additionally, all models demonstrate a strong tendency to associate specific ethnic groups with particular religions. For example, Buddhism is predominantly linked to Asian individuals, reinforcing cultural stereotypes. Such strong associations may inadvertently weaken the public’s perception of the diversity of certain demographic groups and reinforce stereotypes, leading to overly generalized or inaccurate representations of real-world populations.
4 Related Work
Recent studies have focused on both text and multimodal settings, showing that biases persist in various forms, including sexual objectification and cultural conflation Wolfe et al. (2023); Wolfe and Caliskan (2022); Wang et al. (2024b), misclassification tied to skin tone Hazirbas et al. (2024), diagnostic inequities Yang et al. (2025), and a pronounced masculine tilt Huang et al. (2024). BiasAsker Wan et al. (2023), Zhao et al. (2024) and Huang et al. (2025a) reveal significant social biases in LLM outputs. Gupta et al. (2024) show persona-based prompts worsen stereotypes, and Marked Personas Cheng et al. (2023) highlights how explicit demographic references magnify biased portrayals. Fact-or-Fair BiasPainter Wang et al. (2024a) and Ghosh and Caliskan (2023) uncover biased image generation, while Ruggeri and Nozza (2023), Brinkmann et al. (2023), Srinivasan and Bisk (2022), and Mandal et al. (2023) emphasize how model size, objectives, and representations can exacerbate biases.
Existing papers provide datasets and benchmarks for systematically evaluating social biases in AI Ross et al. (2021); Hall et al. (2023); Huang et al. (2025b); Zhou et al. (2022); Du et al. (2025). BIGbench Luo et al. (2024) unifies bias evaluations in text-to-image models. BiasLens Li et al. (2024) offers 33,000 role-specific questions to test fairness in LLMs. Meanwhile, BiasDora Raj et al. (2024) focuses on hidden biases. PAIRS Fraser and Kiritchenko (2024) and Sathe et al. (2024) use synthetic datasets to probe gender, race, and age biases in VLMs. SocialCounterfactuals Howard et al. (2024) provides counterfactual image-text pairs for intersectional bias analysis. Studies also mitigate social bias in VLMs via dataset curation process Hamidieh et al. (2024), an LLM-driven critique and test suite Wan and Chang (2024), counterfactual training data Zhang et al. (2022), and a lightweight post-processing step that curbs gender and racial bias without hurting retrieval accuracy Kong et al. (2023).
While prior studies have explored social biases in VLMs, they often focus exclusively on either explicit or implicit manifestations, or are limited to specific modalities such as text or generated images. In contrast, VisBias evaluates both explicit and implicit biases within a unified experimental design. Notably, our form completion scenario introduces a novel method for probing implicit bias by asking VLMs to complete a 20-attribute personal information form. It is the first to jointly assess model outputs in direct judgment and assistance contexts, providing a holistic view of bias in state-of-the-art VLMs.
5 Conclusions
This study uses VisBias to evaluate five popular VLMs, revealing relatively modest explicit bias in direct multiple-choice or yes-no scenarios, and more subtle, implicit biases persist—most prominently in tasks like image description and form completion. These tasks reveal troubling patterns, such as rigid associations between demographic attributes (e.g., certain ethnicities and specific religions) and gender-linked stereotypes in occupational or personality inferences. We emphasize the need for mitigation strategies that address not only conspicuous, intentional discrimination but also the subtler, unconscious biases in VLM outputs.
Limitations
This study has several limitations. (1) U.S.-centric demographic framing: We adopt the five racial labels Asian, Black, Hispanic, Middle Eastern, and White and rely on occupational imbalance statistics from the U.S. Bureau of Labor Statistics. These choices align with widely used U.S. taxonomies and facilitated data collection, yet they inevitably exclude many identities (e.g., Indigenous peoples, multiracial categories) and occupational patterns that prevail outside the U.S. (2) Selection of question options and attributes: In the explicit MCQ scenario we supply six income brackets, four education levels, four political orientations and four religions. By contrast, the implicit form completion task includes twenty attributes, eight education levels and six religions (including “None”). These design decisions were guided by (i) coverage in prior VLM bias work and (ii) the need to keep MCQs short enough to fit within model context limits. (3) The division of explicit and implicit biases: The distinction between MCQ/Yes-No tasks and image description/form completion reflects a pragmatic design choice. However, the evaluation of explicit and implicit biases is not confined to these tasks. For instance, MCQ can also be leveraged to revel implicit biases.
Ethics Statements
This work investigates social biases in VLMs using publicly available images and model outputs. To elicit potential biases, we applied Caesar cipher-based prompts; these are not intended to generate harmful outputs but to surface otherwise hidden model behaviors. Human annotations are collected via Prolific with fair compensation and anonymized responses. Examples of biased outputs are included for transparency and research purposes only.
Appendix A Experiment Overview
Model MCQ w/o J MCQ w/ J Yes-No w/o J Yes-No w/ J Description Form
GPT-4V Refuse (99.7)
GPT-4o Refuse (100) Refuse (100)
Gemini-1.5-Pro Refuse (98.8)
LLaMA-3.2-Vision Unable (100) Unable (100) Refuse (94.3)
LLaVA-v1.6-13B Unable (98.1) Unable (57.3)
Midjourney N/A N/A N/A N/A N/A
Appendix B Full Result Tables